# Agnostic Smoothed Online Learning

### v2

@@ 4.1 A simple algorithm for a weaker regret guarantee
since the functions $f_{k}$ and $f^{\star}$ agreed on all queries of previous epochs. We can then use Proposition 9 which bounds the sum for each epoch $k\in[K]$. Applying Proposition 9 for $q\geq q_{0}:=\ln^{2}(T)/(K\sigma)$ bounds the number of epochs for which this sum deviates significantly. At the high level, it implies that the quantities $\gamma_{T_{k-1}}(t)$ are roughly of order $q_{0}$ in average. After computations, we obtain
$\mathbb{E}\left[\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)\right]\lesssim\frac{\ln^{3}T}{K\sigma}\cdot T.$ (4)
Putting the two regret terms from Eqs. 3 and 4 together and optimizing over the choice of $K$ gives the same bound as Theorem 8 for the expected oblivious regret of Cover. We give some ideas about how this oblivious regret guarantee can be turned into an adaptive regret guarantee in Section 4.3.
@@ 4.2 Achieving the optimal regret using recursive covers
We use a localization strategy to increase the number of effective epochs on which a cover is recomputed. To not incur a large regret term due to the learning with expert algorithms, we introduce an adaptive variant from the classical Hedge algorithm, A-Exp, which has a regret bound depending on the some notion of difficulty of the learning with expert problem instead of a worst-case bound (see Section 3.3 for a detailed exposition). Going back to an epoch $(T_{k-1},T_{k}]$ of Cover, Proposition 9 essentially implies that during most epochs $k\in[K]$ one can bound
@@
with reasonable probability $1-\delta$. Here, $\hat{y}_{t}$ denotes the predictions of the learning with expert advice algorithm, and we omitted lower-order terms which may depend on the probability failure $\delta$. The regret obtained should be compared to a more classical worst-case bound of order $\sqrt{(T_{k-1}-T_{k})\ln|S|}$ for the Hedge algorithm.
@@
To achieve the optimal regret up to logarithmic terms, we run this strategy recursively over $\left\lfloor\log_{2}(T)/2\right\rfloor$ depths, which is R-Cover. This strategy is akin to some form of chaining at the algorithmic level. The smallest sub-epochs on the last layer have length of order $\sqrt{T}$. Note that the labeled dataset $S$ that is used as parameter in the recursive construction of R-Cover in Algorithm 1 now corresponds to the possible labelings of queries in prior epochs. In practice, the optimal depth to achieve the correct regret dependency in the adversary smoothness parameter $\sigma$ depends on $\sigma$ itself. To avoid requiring this information when implementing R-Cover, at each depth, in addition to the experts corresponding to the predictions of the next layer algorithm, we also add an expert that uses a single function as prediction ($f_{S}$ in line 7 of Algorithm 1). The rationale is that this hedges the final algorithm for all choices of depths at once. Within the proof, we may then focus on the algorithm up to a fixed $\sigma$-dependent depth.
@@ 4.3 Proof sketch for Theorem 4
Now that we have introduced the main conceptual ingredients of the proof, we give a brief sketch of the regret bound. We start by focusing on the oblivious regret compared to some fixed benchmark function $f^{\star}$. R-Cover is composed of $P$ layers. Each layer $p\in[P]$ corresponds to epochs $(T_{k-1}^{(p)},T_{k}^{(p)}]$ for $k\in[N_{p}]$. For instance, initially there is a single epoch for $p=P$ and at the last layer $p=0$ there are $\approx\sqrt{T}$ epochs. For each depth $p\in[P]$ and epoch $k\in[N_{p}]$ we consider the depth-$p$ R-Cover algorithm that was instantiated with the “correct” labeling according to $f^{\star}$. That is, we focus on the algorithm that used the dataset
$S_{k}^{(p)}=\left\{(x_{t},f^{\star}(x_{t})),t\in[T_{k-1}^{(p)}]\right\}.$
@@
The main point of the recursive procedure is that it allows to localize the error by focusing only on the runs of R-Cover that used these correct labeled datasets. The first step of the proof (Section 5.2) is to show that we can decompose the regret of the algorithm compared to $f^{\star}$ in the following way, where $\hat{y}_{t}$ denotes the predictions of the final algorithm. With probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))\lesssim\sum_{p=p_{0}}^{P}\sum_{k\in[N_{p}]}\underbrace{\sqrt{\max\left(\Delta_{k}^{(p)},2\right)d\ln T}}_{Reg_{k}^{(p)}}\\ +\sum_{k\in[N_{p_{0}}]}\underbrace{\sum_{t=T_{k-1}^{(p_{0})}}^{T_{k}^{(p_{0})}}\ell_{t}\left(f_{k,S}^{(p_{0})}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))}_{\Lambda_{k}^{(p_{0})}}+N_{p_{0}}\ln\frac{1}{\delta}.$ (6)
The first term of Eq. 6 corresponds to the regret accumulated along the localization trajectory for running the learning with expert advice algorithm A-Exp. Up to minor details, here $Reg_{k}^{(p)}$ corresponds to the bound on the regret incurred by A-Exp for the depth-$p$ algorithm run during epoch $k\in[N_{p}]$, which is guaranteed by Lemma 7. The quantity $\Delta_{k}^{(p)}$ is the same as that which appears in Lemma 7 and measures the difficulty of the learning with expert problem on epoch $k$ at depth $p$. Technically, the bound from Lemma 7 also includes a failure probability term which accumulated over the complete trajectory corresponds to the term $N_{p_{0}}\ln\frac{1}{\delta}$. This can be viewed as a lower order term.
The second term of Eq. 6 intuitively corresponds to the excess error of a learner that at the beginning of each depth-$p_{0}$ epoch $k\in[N_{p_{0}}]$ has access to an oracle which reveals the values of the optimal function $f^{\star}$ on all prior epoch queries $(x_{t})_{t\leq T_{k-1}^{(p_{0})}}$. Here, we use the notation $f_{k,S}^{(p_{0})}$ to denote the base function $f_{S}$ that was constructed either line 2 or 4 of Algorithm 1 during the run of the depth-$p_{0}$ algorithm R-Cover on epoch $k$ using the correct labeling $S_{k}^{(p_{0})}$. The quantity $\Lambda_{k}^{(p_{0})}$ corresponds to the excess error of this base function $f_{k,S}^{(p_{0})}$ compared to $f^{\star}$ on the epoch $k$.
@@
The next step of the proof is to bound each regret term and more precisely to bound the terms $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$. Using the same arguments as described in Section 4.1 when bounding the excess error of Cover on each epoch, we can show that the quantity
$\Gamma_{k}^{(p)}:=\sum_{t=T_{k-1}^{(p)}+1}^{T_{k}^{(p)}}\gamma_{T_{k-1}^{(p)}}(t)$
bounds both $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$ up to constant factors. Within the general proof for regression, these need to be bounded by different quantities $\Gamma_{k}^{(p,2)}$ and $\Gamma_{k}^{(p,1)}$ respectively in which $\gamma(t)$ is replaced by the $\ell_{2}$ and $\ell_{1}$ deviations from $f^{\star}$ (see Lemma 12 for the precise bound). In classification, because values lie in $\{0,1\}$ both norms are identical hence we can use a single quantity $\Gamma_{k}^{(p)}$.
The terms $\Gamma_{k}^{(p)}$ can now be bounded using Proposition 9, which is the only step so far that required the smoothness assumption on the adversary. For regression, the generalization of this result is Lemma 13. The proof of this main lemma is given in Section 5.5. In this full form, the guarantee obtained on $\Gamma_{k}^{(p,1)}$ and $\Gamma_{k}^{(p,2)}$ depends on the scale $\epsilon$ of the cover constructed at the beginning of each epoch (in classification we used $\epsilon=0$ in Algorithm 1). Naturally, the guarantee degrades as $\epsilon$ grows.
@@
We first construct a cover of the function class $\mathcal{F}$ with respect to the base measure $\mu$ and aim to have low regret compared to functions in this cover. Precisely, we construct a subset of $\mathcal{F}$ such that for all $f\in\mathcal{F}$ there exists $h\in\mathcal{H}$ with
$\mathbb{P}_{x\sim\mu}(f(x)\neq h(x))\leq\epsilon.$
Since $\mathcal{F}$ has VC dimension $d$, we can ensure $\ln|\mathcal{H}|\leq 2d\ln(e^{2}/\epsilon)$ (see [32] or [12, Lemma 13.6]). Using the union bound over the high-probability oblivious regret bounds from the previous section, we can ensure that R-Cover has low regret compared to all functions within the cover. We note that contrary to [36], this covering construction is only for proof purposes and is not performed within the algorithm R-Cover. In fact, since $\mu$ is unknown in our setting, computing such a cover is impossible, as exemplified by the lower bound Theorem 5.
It then only remains to show that $\mathcal{H}$ is also a good cover on the queries made by the smooth adversary $(x_{t})_{t\in[T]}$. Note that this would be immediate if these queries are i.i.d. sampled from $\mu$ from standard VC uniform convergence bounds. To reduce to the i.i.d. case, [36] show the following coupling lemma.
@@
Let $(X_{t})_{t\in[T]}$ be $\sigma$-smooth with respect to $\mu$. Then for all $k\geq 1$, there exists a coupling of $(X_{t})_{t\in[T]}$ with random variables $\{Z_{t,j},t\in[T],j\in[k]\}$ such that the $Z_{t,j}\overset{iid}{\sim}\mu$ and on an event $\mathcal{E}_{k}$ of probability at least $1-Te^{-\sigma k}$, we have $X_{t}\in\{Z_{t,j},j\in[k]\}$ for all $t\in[T]$.
@@ 5 Regret analysis
In this section, we first describe our full algorithm R-Cover for regression in Section 5.1 then prove our main results Theorems 4 and 6 in the rest of the section.
@@ 5.1 General recursive procedure for regression
In this section, we generalize the algorithm given for classification in Algorithm 1 to handle general regression function classes $\mathcal{F}$. Note that at the beginning of each new epoch, R-Cover effectively computes a 00-cover of the previously observed dataset. In the regression setting, we instead compute an $\epsilon$-cover for some $\epsilon>0$ of the functions within the class $\mathcal{F}$ centered around a reference function $f_{0}\in\mathcal{F}$. This effectively restricts the search space of the algorithm to the neighborhood of $f_{0}$, and replaces the labeled dataset $S$ from Algorithm 1 in the classification case (these will be equivalent in this case). Instead of using a single reference function $f_{0}$, we use a sequence of reference functions which will correspond to reference functions from previous depths. This is used to ensure that the search space within $\mathcal{F}$ of sub-algorithms (akin to line 7 of Algorithm 1) are consistent with the search space of algorithm calls from previous depths. These reference functions $f_{i}:\mathcal{X}\to[0,1]$ are stored together with the start time of their corresponding algorithm call $t_{i}\in[T]$, within a set $S$.
@@ 5.2 Regret decomposition
Let us now focus on a single depth-$p$ epoch $k\in[N_{p}]$ for $p>0$. This is composed of $2$ sub-epochs $E_{2(k-1)+l}^{(p-1)}$ for $l\in[2]$. On each sub-epoch $l\in[2]$, the algorithm performs the exponentially weighted algorithm using experts that we denote $\mathcal{A}_{l,r^{\prime}}^{(p-1)}$ for $r^{\prime}\in[r_{l}^{(p-1)}]$. We also denote by $\mathcal{A}_{l,r^{\prime}}^{(p-1)}(t)$ their prediction for some time during the corresponding epoch $E^{(p-1)}_{2(k-1)+l}$. Last, we use the following notation to denote the magnitude of the expert problem at epoch $l$, where $\hat{p}_{t}$ denotes the distribution over the experts $\mathcal{A}_{l,r^{\prime}}^{(p-1)}(t)$ that was used by A-Exp at time $t$:
@@
where by abuse of notation we kept $R_{r^{\prime},t}$ to denote the cumulative regret compared to algorithm $r^{\prime}$ up to time $t$ during epoch $E_{2(k-1)+l}^{(p-1)}$. For convenience, let us denote by $Reg_{k,l}^{(p)}$ the regret incurred by the exponentially weighted algorithm A-Exp on each epoch $E_{2(k-1)+l}^{(p-1)}$ for $l\in[2]$ (see line 7 of Algorithm 4). Then,

### v3

@@ 4.1 A simple algorithm for a weaker regret guarantee
since the functions $f_{k}$ and $f^{\star}$ agreed on all queries of previous epochs. We can then use Proposition 7 which bounds the sum for each epoch $k\in[K]$. Applying Proposition 7 for $q\geq q_{0}:=\ln^{2}(T)/(K\sigma)$ bounds the number of epochs for which this sum deviates significantly. At the high level, it implies that the quantities $\gamma_{T_{k-1}}(t)$ are roughly of order $q_{0}$ in average. Precisely, letting $\Delta T:=\max_{k\in[K]}T_{k}-T_{k-1}=\mathcal{O}(T/k)$ and $l_{0}:=\left\lfloor\log_{2}(1/q_{0})\right\rfloor=\mathcal{O}(\log T)$, Proposition 7 implies that with probability at least $1-(l_{0}+1)\delta$,
$\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)$ $\leq q_{0}T+\sum_{l=0}^{l_{0}}\sum_{k\in[K]}\overbrace{\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)\cdot\mathbbm{1}[\gamma_{T_{k-1}}(t)\in[2^{l}q_{0},2^{l+1}q_{0}]]}^{\Gamma_{k,l}}$
$\leq q_{0}T+(l_{0}+1)K\cdot w(T,\delta)+\sum_{l=0}^{l_{0}}2^{l+1}q_{0}\Delta T\cdot\left|\left\{k\in[K]:\Gamma_{k,l}\geq w(T,\delta)\right\}\right|$
$\leq q_{0}T+(l_{0}+1)K\cdot w(T,\delta)+\sum_{l=0}^{l_{0}}2^{l+1}q_{0}\Delta T\cdot C\frac{\ln^{2}T}{2^{l}q_{0}\sigma}\lesssim\frac{\ln^{3}T}{K\sigma}\cdot T.$ (5)
Putting the two regret terms from Eqs. 4 and 4.1 together and optimizing over the choice of $K$ gives the same bound as Theorem 6 for the expected oblivious regret of Cover. We explain how this oblivious regret guarantee can be turned into an adaptive regret guarantee in Section 4.3.
@@ 4.2 Achieving the optimal regret using recursive covers
We use a localization strategy to increase the number of effective epochs on which a cover is recomputed. To not incur a large regret term due to the learning with expert algorithms, we introduce the adaptive variant from the classical Hedge algorithm, A-Exp, which has a regret bound depending on the actual difficulty of the learning with expert problem instead of a worst-case bound (see Section 3.2). Going back to an epoch $(T_{k-1},T_{k}]$ of Cover, Proposition 7 essentially implies that during most epochs $k\in[K]$ one can bound
@@
with reasonable probability $1-\delta$, instead of the worst-case bound $\sqrt{(T_{k-1}-T_{k})\ln|S|}$ for the Hedge algorithm. Here, $\hat{y}_{t}$ denotes the predictions of the learning with expert advice algorithm, and we omitted lower-order terms which may depend on the probability failure $\delta$.
@@
To achieve the optimal regret, we run this strategy recursively over $\left\lfloor\log_{2}(T)/2\right\rfloor$ depths, which is R-Cover. This strategy is akin to some form of chaining at the algorithmic level. The smallest sub-epochs on the last layer have length of order $\sqrt{T}$. Note that the labeled dataset $S$ that is used as parameter in the recursive construction of R-Cover in Algorithm 1 now corresponds to the possible labelings of queries in prior epochs. In practice, the optimal depth to achieve the best regret regret bound depends depends on the smoothness parameter $\sigma$. To avoid requiring this information when implementing R-Cover, at each depth, in addition to the experts corresponding to the predictions of the next layer algorithm, we also add an expert that uses a single function as prediction ($f_{S}$ in lines 6 and 10 of Algorithm 1). This hedges the final algorithm for all choices of depths at once. Within the proof, we may then focus on the algorithm up to a fixed $\sigma$-dependent depth.
@@ 4.3 Proof sketch for Theorem 2
We start by focusing on the oblivious regret compared to some fixed benchmark function $f^{\star}$. R-Cover is composed of $P$ layers. Each layer $p\in[P]$ corresponds to epochs $(T_{k-1}^{(p)},T_{k}^{(p)}]$ for $k\in[N_{p}]$. For instance, initially there is a single epoch for $p=P$ and at the last layer $p=0$ there are $\approx\sqrt{T}$ epochs. For each depth $p\in[P]$ and epoch $k\in[N_{p}]$ we consider the depth-$p$ R-Cover algorithm that was instantiated with the “correct” labeling according to $f^{\star}$. That is, we focus on the algorithm that used the dataset $S_{k}^{(p)}=\left\{(x_{t},f^{\star}(x_{t})),t\in[T_{k-1}^{(p)}]\right\}$.
@@
The point of the recursive procedure is that it allows to localize the error by focusing only on the runs of R-Cover that used these correct labeled datasets. The first step of the proof (Section 5.2) is to show that we can decompose the regret of the algorithm compared to $f^{\star}$ in the following way, where $\hat{y}_{t}$ denotes the predictions of the final algorithm. With probability $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))\\ \lesssim\sum_{p=p_{0}}^{P}\sum_{k\in[N_{p}]}\underbrace{\sqrt{\max\left(\Delta_{k}^{(p)},2\right)d\ln T}}_{Reg_{k}^{(p)}}+\sum_{k\in[N_{p_{0}}]}\underbrace{\sum_{t=T_{k-1}^{(p_{0})}}^{T_{k}^{(p_{0})}}\ell_{t}\left(f_{k,S}^{(p_{0})}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))}_{\Lambda_{k}^{(p_{0})}}+N_{p_{0}}\ln\frac{1}{\delta}.$ (7)
The first term of Eq. 7 is the regret accumulated along the localization trajectory for running the learning with expert advice algorithm A-Exp. Up to minor details, here $Reg_{k}^{(p)}$ corresponds to the bound on the regret incurred by A-Exp for the depth-$p$ algorithm run during epoch $k\in[N_{p}]$, which is guaranteed by Lemma 5. The quantity $\Delta_{k}^{(p)}$ is the same as that which appears in Lemma 5 and measures the difficulty of the learning with expert problem on epoch $k$ at depth $p$. Technically, the bound from Lemma 5 also includes a failure probability term which accumulated over the complete trajectory corresponds to the term $N_{p_{0}}\ln\frac{1}{\delta}$. This can be viewed as a lower order term.
The second term of Eq. 7 intuitively corresponds to the excess error of a learner that at the beginning of each depth-$p_{0}$ epoch $k\in[N_{p_{0}}]$ has access to an oracle which reveals the values of the optimal function $f^{\star}$ on all prior epoch queries $(x_{t})_{t\leq T_{k-1}^{(p_{0})}}$. Here, we use the notation $f_{k,S}^{(p_{0})}$ to denote the base function $f_{S}$ that was constructed at line 1 of Algorithm 1 during the run of the depth-$p_{0}$ algorithm R-Cover on epoch $k$ using the correct labeling $S_{k}^{(p_{0})}$. The quantity $\Lambda_{k}^{(p_{0})}$ corresponds to the excess error of this base function $f_{k,S}^{(p_{0})}$ compared to $f^{\star}$ on the epoch $k$.
@@
We next bound each regret term and more precisely bound the terms $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$. Using the same arguments as described in Section 4.1 when bounding the excess error of Cover on each epoch, we can show that the quantity
$\Gamma_{k}^{(p)}:=\sum_{t\in(T_{k-1}^{(p)},T_{k}^{(p)}]}\gamma_{T_{k-1}^{(p)}}(t)$
bounds both $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$ up to constant factors. Within the general proof for regression, these need to be bounded by different quantities $\Gamma_{k}^{(p,2)}$ and $\Gamma_{k}^{(p,1)}$ respectively in which $\gamma(t)$ is replaced by the $\ell_{2}$ and $\ell_{1}$ deviations from $f^{\star}$ (see Lemma 10 for the precise bound). We then bound the terms $\Gamma_{k}^{(p)}$ using Proposition 7, which is the only step so far that required the smoothness assumption. For regression, the generalization of this result is Lemma 11. In this full form, the guarantee obtained on $\Gamma_{k}^{(p,1)}$ and $\Gamma_{k}^{(p,2)}$ depends on the scale $\epsilon$ of the cover constructed at the beginning of each epoch (in classification we used $\epsilon=0$ in Algorithm 1). Naturally, the guarantee degrades as $\epsilon$ grows.
@@
We first construct a cover of the function class $\mathcal{F}$ with respect to the base measure $\mu$ and aim to have low regret compared to functions in this cover. Precisely, we construct a subset of $\mathcal{F}$ such that for all $f\in\mathcal{F}$ there exists $h\in\mathcal{H}$ with $\mathbb{P}_{x\sim\mu}(f(x)\neq h(x))\leq\epsilon$. Since $\mathcal{F}$ has VC dimension $d$, we can ensure $\ln|\mathcal{H}|\leq 2d\ln(e^{2}/\epsilon)$ (see [32] or [12, Lemma 13.6]). Using the union bound over the high-probability oblivious regret bounds from the previous section, we can ensure that R-Cover has low regret compared to all functions within the cover. We note that contrary to [36], this covering construction is only for proof purposes and is not performed within the algorithm R-Cover. In fact, since $\mu$ is unknown in our setting, computing such a cover is impossible, as exemplified by the lower bound Theorem 3.
It then only remains to show that $\mathcal{H}$ is also a good cover on the queries made by the smooth adversary $(x_{t})_{t\in[T]}$. This would be immediate if these queries are i.i.d. sampled from $\mu$ by VC uniform convergence. To reduce to the i.i.d. case, [36] show the following coupling lemma.
@@
Let $(X_{t})_{t\in[T]}$ be $\sigma$-smooth with respect to $\mu$. Then for all $k\geq 1$, there is a coupling of $(X_{t})_{t\in[T]}$ with random variables $\{Z_{t,j},t\in[T],j\in[k]\}$ such that the $Z_{t,j}\overset{iid}{\sim}\mu$ and on an event $\mathcal{E}_{k}$ of probability at least $1-Te^{-\sigma k}$, we have $X_{t}\in\{Z_{t,j},j\in[k]\}$ for all $t\in[T]$.
@@ 5 Regret analysis
In this section, we first describe our full algorithm R-Cover for regression in Section 5.1 then prove our main results Theorems 2 and 4 in the rest of the section.
@@ 5.1 General recursive procedure for regression
In this section, we generalize the algorithm given for classification in Algorithm 1 to handle general regression function classes $\mathcal{F}$. Note that at the beginning of each new epoch, R-Cover effectively computes a 00-cover of the previously observed dataset. In the regression setting, we instead compute an $\epsilon$-cover for some $\epsilon>0$ of the functions within the class $\mathcal{F}$ centered around a reference function $f_{0}\in\mathcal{F}$. This effectively restricts the search space of the algorithm to the neighborhood of $f_{0}$, and replaces the labeled dataset $S$ from Algorithm 1 in the classification case (these will be equivalent in this case). Instead of using a single reference function $f_{0}$, we use a sequence of reference functions which will correspond to reference functions from previous depths. This is used to ensure that the search space within $\mathcal{F}$ of sub-algorithms (akin to lines 6 and 10 of Algorithm 1) are consistent with the search space of algorithm calls from previous depths. These reference functions $f_{i}:\mathcal{X}\to[0,1]$ are stored together with the start time of their corresponding algorithm call $t_{i}\in[T]$, within a set $S$.
@@ 5.2 Regret decomposition
Let us now focus on a single depth-$p$ epoch $k\in[N_{p}]$ for $p>0$. This is composed of $2$ sub-epochs $E_{2(k-1)+l}^{(p-1)}$ for $l\in[2]$. On each sub-epoch $l\in[2]$, the algorithm performs the exponential weights algorithm using experts that we denote $\mathcal{A}_{l,r^{\prime}}^{(p-1)}$ for $r^{\prime}\in[r_{l}^{(p-1)}]$. We also denote by $\mathcal{A}_{l,r^{\prime}}^{(p-1)}(t)$ their prediction for some time during the corresponding epoch $E^{(p-1)}_{2(k-1)+l}$. Last, we use the following notation to denote the magnitude of the expert problem at epoch $l$, where $\hat{p}_{t}$ denotes the distribution over the experts $\mathcal{A}_{l,r^{\prime}}^{(p-1)}(t)$ that was used by A-Exp at time $t$:
@@
where by abuse of notation we kept $R_{r^{\prime},t}$ to denote the cumulative regret compared to algorithm $r^{\prime}$ up to time $t$ during epoch $E_{2(k-1)+l}^{(p-1)}$. For convenience, let us denote by $Reg_{k,l}^{(p)}$ the regret incurred by the exponential weights algorithm A-Exp on each epoch $E_{2(k-1)+l}^{(p-1)}$ for $l\in[2]$ (see line 7 of Algorithm 4). Then,